OpenStudy (anonymous):

Find the quotient 3^10/3^-2

ganeshie8 (ganeshie8):

hint : \(\large \dfrac{a^m}{a^n} = a^{m-n}\)

OpenStudy (anonymous):

i dont have a good calculator so what is 10-(-2)

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

10 - (-2) = 10 + 2 = 12

OpenStudy (anonymous):

TY SO Much

ganeshie8 (ganeshie8):

yw

OpenStudy (anonymous):

What about 10^2/10^5 @ganeshie8 Would that be 10^-3